OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of the function. limit of x divided by the quantity x minus 5 as x approaches 5 from the left

OpenStudy (anonymous):

@hoblos @undeadknight26

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

Is this your question \[ \lim_{x\to 5^{-}} \frac x{x-5} \]

OpenStudy (anonymous):

x is >0 near 5 and x-5 <0 when x approaches 5 from the left Hence \[ \lim_{x\to 5^{-}} \frac x{x-5}=-\infty \] The graph above confirms that. The line x=5 is vertical asymptote